OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge \frac{\text d}{\text dx} x^{e^{x^2}}\]

OpenStudy (anonymous):

rewrite as \[e^{e^{x^2}\ln(x)}\] and use the chain rule

OpenStudy (lgbasallote):

do you mean \[\ln y = e^{x^2} \ln x?\]

OpenStudy (anonymous):

\[\huge x^{e^{x^2}}*lnx* e^{x^2}*2x\]

OpenStudy (lgbasallote):

uhh how?

OpenStudy (anonymous):

\[\huge y=a^{f(x)}\] \[\huge y'=a^{f(x)}*f'(x)*lna\]

OpenStudy (lgbasallote):

oh i see

OpenStudy (lgbasallote):

yeah that makes sense thanks

OpenStudy (anonymous):

I made a mistake..

OpenStudy (anonymous):

x is not constant that formula is for constants

OpenStudy (anonymous):

\[\huge x^{e^{x^2}}*\left( e^{x^2}*2x*lnx+\frac{e^{x2}}{x}\right)\]

OpenStudy (anonymous):

@mukushla you're right, but still we need that formula..

OpenStudy (anonymous):